Accelerator Update for February 24, 2006 - February 27, 2006
Friday February 24
The day shift began with the Tevatron (TeV) cryo system warming up, with Operations monitoring the Recycler stash, with the Antiproton Source (Pbar) stacking, with D-Zero (D0) and the Collider Detector at Fermilab (CDF) in access, and with NuMI, MiniBooNE, and the Meson MCenter experiments taking beam.
Recycler and Pbar experts were conducting studies by 8 AM.
MTest resumed taking beam at 9:48 AM.
Saturday February 25
The midnight shift began with Recycler studies, with Pbar off, and with NuMI, MiniBooNE, MTest, and MCenter taking beam.
At 9:37 AM, a Booster RF station (BRF7) tripped off. Experts turned off a power supply (EAPS) that later wouldn’t turn on due to a water flow problem. Experts bypassed the RF station for repairs. They reported at 1:22 PM that they were able to turn on EAPS by recalibrating a flow meter. Booster beam returned at 1:22 PM.
At 5:05 PM, the NuMI beam permit tripped off due to a vacuum valve problem. Operators found no vacuum problem, but when they cycled the valve it closed and wouldn’t open. Operations called in an expert.
At 6:53 PM, a Main Injector (MI) RF station (MIRF16) tripped off and wouldn’t reset. Operators bypassed the station.
Operators prepared the Recycler for shutdown at 11:21 PM.
Sunday February 26
The midnight shift began with the TeV cryo systems warming up, with Pbar off, and with NuMI, MiniBooNE, MTest, and MCenter taking beam.
At 12:03 AM, Operations turned off the Booster, MI, and Linac.
By 12:13 AM, Operations had NuMI and MiniBooNE prepared for access.
By 8 AM, all machines were being prepared for the shutdown to begin.
Monday February 27
The midnight shift began with Fermi Accelerator complex in shutdown for twelve weeks.
The Plans for Monday and the ShutdownThe plans are to begin upgrades, install new equipment, and conduct repairs.
Comments and Suggestions
|last modified 02/27/2006 email Fermilab|